OpenStudy (anonymous):

Suppose JK bisects angle LJM. The measure of angle LJK = (-10x + 3) and the measure of KJM = (-x + 21)º. Find the measure of angle LJM. Need help on how to solve

OpenStudy (mathlegend):

So if I were you I would set these two equations equal to one another. Being that they are said to be bisecting one another. Basically, cutting one another in half.

OpenStudy (mathlegend):

-10x+3 = -x+21

OpenStudy (mathlegend):

Can you solve that equation?

OpenStudy (anonymous):

yes

OpenStudy (mathlegend):

Okay tell me what you get for 'x'

OpenStudy (anonymous):

it's not coming out right

OpenStudy (mathlegend):

Try to keep the 'x' positive. -10x+3=-x+21 +10x +10x ------------ 3=9x+21

OpenStudy (mathlegend):

Can you solve from there?

OpenStudy (anonymous):

yea i can

OpenStudy (anonymous):

i thought you were suppose to add x on both sides

OpenStudy (mathlegend):

Well I like to keep the 'x' positive. I hate negative numbers. lol

OpenStudy (anonymous):

well i know its wrong

OpenStudy (mathlegend):

-10x+3=-x+21 +10x +10x ------------ 3=9x+21 -21 -21 ----------- -18 = 9x -2 = x

OpenStudy (mathlegend):

Do you follow?

OpenStudy (mathlegend):

If you added 'x' to both sides it would come out to the same outcome.

OpenStudy (mathlegend):

You probably made a minor mistake.

OpenStudy (mathlegend):

Now, we must plug this negative 2 into both LJK and KJM

OpenStudy (mathlegend):

-10(-2)+3 = ? -(-2)+21 = ?

OpenStudy (mathlegend):

Solve those @raevenbennett

OpenStudy (anonymous):

i did i got the answer of 23

OpenStudy (mathlegend):

Good so now we know each individual angle equals 23. So we have... LJK = 23 KJM = 23 The entire angle which is: LJM = 46 degrees

OpenStudy (mathlegend):